# Towards Variable-Length In-Network Caching

### v2

@@ Abstract
We present StarCache, a new in-network caching architecture that can cache variable-length items to balance a wide range of key-value workloads. Unlike existing works, StarCache does not cache hot items in the switch memory. Instead, we make hot items revisit the switch data plane continuously by exploiting packet recirculation. Our approach keeps cached key-value pairs in the switch data plane while freeing them from item size limitations caused by hardware constraints. We implement a StarCache prototype on an Intel Tofino switch. Our experimental results show that StarCache can balance highly skewed workloads with various key and value sizes.
@@ 1 Introduction
Key-value stores are the fundamental building blocks for online services owing to fast access to variable-length data [4, 17, 5, 31]. Workloads are generally read-intensive, and most items are hundreds of bytes [13, 8, 9]. A challenge in scaling key-value stores is mitigating load imbalance among storage servers caused by skewed key popularity (e.g., trending events). The load imbalance leads to the overloading of hot item servers, causing significant performance degradation [22, 30, 28].
@@
In-network caching is a promising load-balancing approach that builds a cache in the switch by leveraging programmable switch ASICs like Intel Tofino [6]. The idea is to store hot items in the switch memory. Requests read the item value by referring to the cache lookup table where the item key is the table index. Compared to server-based caching, in-network caching provides much higher performance without extra cache nodes or accelerators thanks to a Tbps-scale processing throughput.
@@
Unfortunately, existing solutions [22, 36, 33] can cache only tiny items due to hardware constraints. Specifically, the key size is limited to 16 bytes due to the maximum match-key width of the match-action table. A limited number of match-action stages and a small accessible byte size per stage make it difficult for the value size to exceed 128 bytes. These numbers are far from typical workloads where key and value sizes are still small enough to be stored in a single packet, but larger than the limit [8, 34, 13, 9, 38]. For example, in 54 Twitter workloads [38], most keys are tens of bytes and many values are less than 1024 bytes. However, the existing solutions cannot cache even a single item in 42 out of 54 workloads because the keys and values are generally larger than the limits. Between the other 12 workloads, only 2 workloads have a portion of cacheable items larger than 50%. This trend is similar in Facebook workloads as well [13]. In this context, we ask the following question: how can we cache variable-length items in the programmable switch for balancing a wide range of key-value workloads?
@@
We present StarCache, a new in-network caching architecture that can cache variable-length key-value items in the switch data plane. Our idea is to make hot items keep visiting the switch data plane continuously in the form of cache packets, instead of caching the hot items in the switch memory. We efficiently use packet recirculation, a built-in feature of the programmable switch ASIC. This enables a packet to revisit the switch data plane through an internal loopback port without going outside. The switch can cache variable-length key-value pairs without being limited by the hardware constraints owing to circulating cache packets. For cache serving, the switch maintains small request metadata in the switch memory and forwards the matched cache packet to the client if there is a pending request. To support variable-length keys, we use key hashes for cache lookups and resolve hash collisions at the client by comparing the requested key and the returned key, which is contained in the cache packet.
@@
The key insight behind the idea is that it is hard to overcome the memory access constraint of switch hardware if we stick to the approach of caching items in the switch memory. For example, the idea of using key hashes for variable-length keys is difficult to realize in the existing architecture. This is because the switch does not have enough hardware resources to store keys and values together. In StarCache, cache packets contain both keys and values, allowing to handle hash collisions at the client. Meanwhile, we may recirculate requests multiple times to read large values. However, this limits scalability because the number of in-flight packets in the recirculation port increases proportionately to the number of requests. In addition, the switch has tens of ports but only one recirculation port. In ours, the recirculation port is not a performance bottleneck. This is because 1) requests are never recirculated; 2) the number of in-flight cache packets is small; 3) the switch can process multiple cache packets simultaneously with massive hardware-level parallelism, thus minimizing queueing between cache packets.
@@
Realizing the idea imposes various technical challenges. First, the switch should buffer multiple request metadata for cached items. We design a request table as a circular queue structure using multiple register arrays. Our request table provides isolated data access among different keys, which means that requests for a key do not have to wait until requests for other keys are handled. Second, a cache packet should serve multiple requests once fetched, rather than a single request. To achieve this, the switch clones the cache packet before forwarding it to the client using the packet replication engine (PRE), a special module of the switch ASIC. Next, we should ensure cache coherence. We design an invalidation-based coherence protocol that can fetch new values and reply to the client simultaneously. Lastly, the switch should deal with dynamic workloads where key popularity and value size change over time. We design a switch control plane mechanism that considers dynamic workloads.
@@
We have implemented a StarCache prototype on an Intel Tofino switch. To evaluate StarCache, we build a testbed consisting of commodity servers, which are interconnected via an Intel Tofino switch. Our experimental results show that StarCache provides load balancing for skewed workloads having diverse item sizes. In addition, StarCache is robust to various workload conditions. We also show that StarCache can handle dynamic workloads where key popularity changes over time.
@@
  • We design StarCache, an in-network caching architecture where the switch can cache variable-length key-value items to balance a wide range of workloads.
  • We propose several techniques to address various challenges in designing a custom switch data plane.
  • We implement a StarCache prototype on Intel Tofino switches and show the efficiency and robustness of StarCache through extensive testbed experiments.
@@ 2.1 In-Network Load-Balancing Caches
Balancing key-value stores with a small cache. Today’s online services rely on key-value stores since a user request may generate hundreds of key-value store requests. In distributed key-value stores, balancing imbalanced loads between storage servers caused by different key popularity is a primary challenge. Caching is a powerful technique to address the challenge, which is based on a theoretical foundation that we can balance loads for $N$ servers (or partitions) by caching the $O(N\log N$) hottest items, regardless of the number of items [16]. Caching on commodity servers is a natural option to build the load-balancing cache [17, 16, 30]. However, the performance of a cache node is not sufficient to handle many requests due to the limited packet processing throughput of CPUs.
@@
(b) The proposed StarCache architecture
Figure 1: Comparison of the high-level idea with the NetCache architecture. In StarCache, clients submit requests, and then circulating cache packets read request metadata. Since both keys and values are in cache packets, hardware constraints do not limit the item size. For variable-length keys, StarCache uses key hashes while handling hash collisions at the client.
Why in-network caching? Building a load-balancing cache in the switch by leveraging programmable switch ASICs like Intel Tofino [6] is an attractive approach. Unlike server-based caching, this approach can provide high-performance load balancing without additional nodes. This is because the switch is highly optimized for packet I/O and can process a few billion packets per second, which is an order of magnitude higher than that of the CPU in servers. The switch also provides a low packet processing delay within hundreds of nanoseconds.
Limited cacheable item size. Existing works [22, 33, 36] demonstrate the efficiency of in-network caching, but they limit the cacheable item size. Specifically, they support items of up to 16-byte keys and 128-byte values. In the reconfigurable match table (RMT) switch architecture [11] like Intel Tofino, the switch data plane consists of $n$ match-action stages [6]. Each stage has a static memory and a few ALUs that can perform simple arithmetic operations on $k$ bytes. The existing works store the value of cached items across multiple stages after fragmentation, limiting the maximum value size to $n\times k$ bytes. Unfortunately, $n\times k$ of the current switch is quite small, and the available number of stages for accessing the value is less than $n$ since other non-caching functions also consume match-action stages. The key size is also limited by the maximum match-key width of the match-action table. This is because the existing works use a match-action table to implement the cache lookup table where the item key is the match key.
Why is it not enough? Many key-value items are indeed small, but typically exceed the existing size limits. We have analyzed 54 Twitter workloads [38], and observe that existing solutions are not enough to deal with typical workloads. For example, only 3.7% of the workloads have over 80% of keys $\leq 16$ B. 38.9% of the workloads have over 80% of values $\leq 128$ B. The existing works can cache less than 10% of items for 85% of the workloads because, to be cacheable, both key and value sizes must be within the limits. Furthermore, they cannot cache even a single item for 77.8% of the workloads. Facebook workloads [13] show similar size distributions. The average key size is 27.1 bytes, and the median value size is 235 bytes. These numbers exceed the size limits of the existing architecture.
@@ 2.2 Variable-Length In-Network Caching
Design rationale. We argue that it is hard to overcome the memory access constraint if we stick to the approach of caching data in the switch memory because the constraint is determined at the time of manufacture. Therefore, we propose StarCache, a different approach to in-network caching.
@@
Our high-level idea is to keep cached key-value pairs circulating within the switch data plane through packet recirculation, a built-in feature of the programmable switch that makes the packet revisit the switch data plane. In our approach, cached items read requests rather than the requests read the cached items. Specifically, clients submit requests, and the switch maintains small request metadata like the client IP address. Cache packets, which are reply packets containing the key and the value of cached items, continue to revisit the data plane via an internal recirculation port while actively checking pending requests. If a request for the key is found, the matched cache packet is forwarded to the client. Otherwise, the cache packet is recirculated. In Figure 1, we illustrate the difference between StarCache and the NetCache architecture [22], which is the reference architecture of existing in-network caching solutions [22, 33, 36].
@@
Variable-length keys. Our approach enables us to support variable-length key-value pairs. For variable-length keys, one possible solution within the limited match-key width of the match-action table is to use a fixed-size key hash as the match key. To handle potential hash collisions, we should compare the requested key and the key of the value that has been read. To do this, the switch data plane should also store the item keys alongside values. Unfortunately, it is hard to realize in the existing architecture because there are not enough match-action stages to store both keys and values. StarCache can realize the idea of using key hashes since the switch data plane maintains cached key-value pairs in the form of circulating reply packets. Therefore, the client can get the correct value from the storage server if the requested and returned keys differ.
@@
Variable-length values. For variable-length values, a possible approach in the existing architecture is recirculating requests. Specifically, requests can access the switch memory multiple times using packet recirculation. However, this approach is not scalable because the number of in-flight packets per second in the recirculation port increases in proportion to the number of requests. For example, if every request is recirculated 8 times to read a 1024-byte value, the effective throughput of the recirculation port is reduced to 1/8 of the bandwidth. Unfortunately, a pipeline in the programmable switch has only one internal recirculation port, while there are tens of regular front ports. This means that the recirculation port can limit the performance.
@@
In StarCache, the recirculation port is not a bottleneck. The switch never recirculates requests as they are in the memory until cache packets read them. We recirculate only a small number of cache packets. The time to recirculate and process a cache packet is a few hundred nanoseconds like normal packets. Furthermore, the switch can process multiple packets simultaneously thanks to hardware-level parallelization, which reduces queueing between cache packets.
@@
Trade-off. There is a trade-off of using recirculation to achieve variable-length in-network caching. Although we are free from size limitations, we should reasonably limit the number of cache entries to avoid long request wait times. Otherwise, the switch may not have enough space to store requests due to many pending requests, or the request latency would increase excessively even if it is stored. This is because, although the switch can process many packets simultaneously, the queueing between cache packets may still happen due to many cache packets beyond the capability of switch hardware. Fortunately, caching a small number of hot items is enough to balance skewed workloads [16, 28].
@@
  • The switch should adapt to changes in the workload, such as key popularity.
Figure 2: StarCache architecture.
3 StarCache Design
3.1 StarCache Architecture
Figure 2 shows the overview of the StarCache architecture.
Switch data plane. The switch data plane in StarCache consists of several custom tables and modules as follows.
@@ 3.1 StarCache Architecture
  • Cloning module comprises a few match-action tables for cloning reply packets. Packet cloning is done by the PRE, a special hardware module in the switch ASIC [6]. The switch clones packets with negligible overhead since the switch duplicates only a packet descriptor pointing to the packet instead of the whole packet.
Meanwhile, the switch invokes the custom processing logic only for StarCache packets by referring to reserved L4 ports. We use UDP to handle messages for better latency like existing works [28, 30, 22] while using TCP for top-$k$ item reports in cache updates. For normal packets, the switch only applies the traditional packet forwarding logic.
Switch control plane. The controller in the switch control plane performs cache updates and cache sizing. It evicts the least popular keys and inserts new hot keys based on server-side periodic top-$k$ reports for uncached keys and the switch-side report for cached keys. Note that while the controller handles key insertion, fetching values is done via the switch data plane. The controller also adjusts the cache size dynamically to deal with dynamic workloads. It also handles switch configurations.
Clients and servers. Clients should specify the operation type, the item key to request, and the key hash. Storage servers run a server application that acts as a shim layer that translates StarCache messages to API calls for key-value stores and vice versa. The server returns replies for read and write requests like regular storage servers. However, in cases of write requests for cached items, the server returns a write reply with the item value to fetch the latest value.
Figure 3: StarCache packet format.

### v3

# Pushing the Limits of In-Network Caching for Key-Value Stores

@@ Abstract
We present OrbitCache, a new in-network caching architecture that can cache variable-length items to balance a wide range of key-value workloads. Unlike existing works, OrbitCache does not cache hot items in the switch memory. Instead, we make hot items revisit the switch data plane continuously by exploiting packet recirculation. Our approach keeps cached key-value pairs in the switch data plane while freeing them from item size limitations caused by hardware constraints. We implement an OrbitCache prototype on an Intel Tofino switch. Our experimental results show that OrbitCache can balance highly skewed workloads and is robust to various system conditions.
@@ 1 Introduction
Key-value stores are the fundamental building blocks for online services owing to fast access to variable-length data [3, 16, 4, 30]. Workloads are generally read-intensive, and most items are hundreds of bytes [12, 7, 8]. A challenge in scaling key-value stores is mitigating load imbalance among storage servers caused by skewed key popularity (e.g., trending events). The load imbalance leads to overloading hot item servers, causing overall performance degradation [21, 29, 27].
@@
In-network caching is a promising load-balancing approach that builds a cache in the switch by leveraging programmable switch ASICs like Intel Tofino [5]. The idea is to store hot items in the on-chip switch memory. Requests read the item value by referring to the cache lookup table where the item key is the table index. Compared to server-based caching, in-network caching provides much higher performance without extra cache nodes or accelerators thanks to a Tbps-scale processing throughput.
@@
Unfortunately, existing solutions [21, 34, 32] can cache only tiny items due to hardware constraints. Specifically, the key size is limited to 16 bytes due to the maximum match-key width of the match-action table. A limited number of match-action stages and a small accessible byte size per stage make it difficult for the value size to exceed 128 bytes. These numbers are far from typical workloads where key and value sizes are still small enough to be stored in a single packet, but larger than the limit [7, 33, 12, 8, 37]. For example, in 54 Twitter workloads [37], most keys are tens of bytes and many values are less than 1024 bytes. However, the existing solutions cannot cache even a single item in 42 out of 54 workloads because the keys and values are generally larger than the limits. Between the other 12 workloads, only 2 workloads have a portion of cacheable items larger than 50%. This trend is similar in Facebook workloads as well [12]. In this context, we ask the following question: how can we cache variable-length items in the programmable switch for balancing typical key-value workloads?
@@
We present OrbitCache, a new in-network caching architecture that can cache variable-length key-value items in the switch data plane. Our idea is to make hot items visit the switch data plane continuously in the form of cache packets instead of caching the hot items in the switch memory. We efficiently use packet recirculation, a built-in feature of the programmable switch ASIC. This enables a packet to revisit the switch data plane through an internal loopback port without going outside. The switch can cache variable-length key-value pairs without being limited by the hardware constraints owing to circulating cache packets. For cache serving, the switch maintains small request metadata in the switch memory and forwards the cache packet to the client if there is a pending request for the key. To support variable-length keys, we use key hashes for cache lookups and resolve hash collisions at the client by comparing the requested key and the returned key, which is contained in the cache packet.
@@
The key insight behind the idea is that it is hard to overcome the memory access constraint of switch hardware if we stick to the approach of caching items in the switch memory. For example, the idea of using key hashes for variable-length keys is difficult to realize in the existing solutions. This is because the switch does not have enough hardware resources to store keys and values together. In OrbitCache, cache packets contain both keys and values, allowing to handle hash collisions at the client. Meanwhile, we may recirculate requests multiple times to read larger values. However, this limits scalability because the number of in-flight packets in the recirculation port increases proportionately to the number of requests, making a bottleneck. In addition, the switch has one internal recirculation port. OrbitCache avoids a bottleneck in the recirculation port. This is because 1) requests are never recirculated; 2) the number of in-flight cache packets is small and constant; 3) the switch can process multiple cache packets simultaneously with hardware-level parallelism, thus minimizing queueing between cache packets.
@@
Realizing the idea imposes various technical challenges. First, the switch should buffer multiple request metadata for cached items. We design a request table as a circular queue structure using multiple register arrays. Our request table provides isolated data access among different keys. Second, a cache packet should serve multiple requests once fetched, rather than a single request. To achieve this, the switch clones the cache packet with low overhead before forwarding it to the client using the packet replication engine (PRE), a special module of the switch ASIC. Next, we should ensure cache coherence. We design an invalidation-based coherence protocol that can fetch new values and send back a reply to the client simultaneously. Lastly, the switch should deal with dynamic workloads where key popularity changes over time. We design an efficient cache update mechanism in the switch control plane.
@@
We have implemented a OrbitCache prototype on an Intel Tofino switch. To evaluate OrbitCache, we build a testbed consisting of commodity servers. Our experimental results show that OrbitCache provides load balancing for many skewed workloads having diverse item sizes. In addition, OrbitCache is robust to various workload conditions like key access distributions, write ratio, and the number of servers. We also show that OrbitCache can adapt to dynamic workloads where key popularity changes over time.
@@
  • We design OrbitCache, an in-network caching architecture where the switch can cache variable-length hot items to balance a wide range of workloads for distributed key-value stores.
  • We propose several techniques to address technical challenges when realizing the idea of variable-length in-network caching.
  • We implement a OrbitCache prototype on Intel Tofino switches and show the efficiency and robustness of OrbitCache through extensive testbed experiments.
@@ 2.1 In-Network Load-Balancing Caches
Balancing key-value stores with a small cache. In distributed key-value stores, balancing imbalanced loads across storage servers caused by different key popularity is a primary challenge. Caching is a powerful technique to address the challenge, which is based on a theoretical result called the small cache effect: we can balance loads for $N$ servers (or partitions) by caching the $O(N\log N$) hottest items, regardless of the number of items [15]. Caching on commodity servers is a natural option to build the cache [16, 15, 29]. However, the performance of a cache node is not sufficient to handle many requests due to the limited throughput of CPUs. Building a high-performance caching layer using multiple replicated cache nodes is expensive and causes poor write performance for cache coherence.
@@
(b) The proposed OrbitCache architecture
Figure 1: Comparison of the high-level idea with the NetCache architecture. In OrbitCache, clients submit requests, and then circulating cache packets read request metadata. Since both keys and values are in cache packets, hardware constraints do not limit the item size. For variable-length keys, OrbitCache uses key hashes while handling hash collisions at the client.
Why in-network caching? Building a load-balancing cache in the switch by leveraging programmable switch ASICs like Intel Tofino [5] is an attractive approach. Unlike server-based caching, this approach can provide high-performance load balancing in a single box without additional nodes. This is because the switch is highly optimized for packet I/O and can process a few billion packets per second, which is an order of magnitude higher than that of the CPU in servers. The switch also provides a low packet processing delay within hundreds of nanoseconds.
Limited cacheable item size. Existing works [21, 32, 34] demonstrate the efficiency of in-network caching, but they limit the cacheable item size. Specifically, they support items of up to 16-byte keys and 128-byte values. In the reconfigurable match table (RMT) switch architecture [10] like Intel Tofino, the switch data plane consists of $n$ match-action stages [5]. Each stage has a static memory and a few ALUs that can perform simple arithmetic operations on $k$ bytes. The existing works store the value of cached items across multiple stages after fragmentation, limiting the maximum value size to $n\times k$ bytes. Unfortunately, $n\times k$ of the current switch is quite small, and the available number of stages for accessing the value is less than $n$ since other non-caching functions also consume match-action stages. The key size is also limited by the maximum match-key width of the match-action table in a single match-action stage. This is because the existing works use a match-action table to implement the cache lookup table where the item key is the match key.
Why is it not enough? Many key-value items are indeed small, but typically exceed the existing size limits. We have analyzed 54 Twitter workloads [37], and observe that existing solutions are not enough to deal with typical workloads. For example, only 3.7% of the workloads have over 80% of keys $\leq 16$ B. 38.9% of the workloads have over 80% of values $\leq 128$ B. The existing works can cache less than 10% of items for 85% of the workloads because, to be cacheable, both key and value sizes must be within the limits. Furthermore, they cannot cache even a single item for 77.8% of the workloads. Facebook workloads [12] show similar size distributions. The average key size is 27.1 bytes, and the median value size is 235 bytes. These numbers exceed the size limits of the existing solutions.
@@ 2.2 Variable-Length In-Network Caching
Design rationale. We argue that it is hard to overcome the memory access constraint if we stick to the approach of caching data in the switch memory because the constraint is determined at the time of manufacture. Therefore, we propose OrbitCache, a different approach to in-network caching.
@@
Our high-level idea is to keep cached key-value pairs circulating within the switch data plane through packet recirculation, a built-in feature of the programmable switch that makes the packet revisit the switch data plane. In our approach, cached items read requests rather than the requests read the cached items. Specifically, clients submit requests, and the switch maintains small request metadata like the client IP address. Cache packets, which are reply packets containing the key and the value of cached items, continue to revisit the data plane via a recirculation port while actively checking pending requests. Note that cache packets do not go outside the switch. If a request for the key is found, the matched cache packet is forwarded to the client. Otherwise, the cache packet is recirculated. In Figure 1, we illustrate the difference between OrbitCache and the NetCache architecture [21], which is the reference architecture of existing in-network caching solutions [21, 32, 34].
@@
Variable-length keys. Our approach enables us to support variable-length key-value pairs. For variable-length keys, one possible solution within the limited match-key width of the match-action table is to use a fixed-size key hash as the match key. To handle potential hash collisions, we should compare the requested key and the key of the value that has been read. To do this, the switch data plane should also store the item keys alongside values. Unfortunately, it is hard to realize in the existing architecture because there are not enough match-action stages to store both keys and values. OrbitCache can realize the idea of using key hashes since the switch maintains cached key-value pairs in the form of circulating reply packets. Therefore, the client can get the correct value from the storage server if the requested and returned keys differ.
@@
Variable-length values. For variable-length values, a possible approach in the existing architecture is recirculating requests. Specifically, requests can access the switch memory multiple times using packet recirculation. However, this approach is not scalable because the number of in-flight packets per second in the recirculation port increases in proportion to the number of requests. For example, if every request is recirculated 8 times to read a 1024-byte value, the effective throughput of the recirculation port is reduced to 1/8 of the bandwidth. Unfortunately, a pipeline in the programmable switch has only one internal recirculation port, while there are tens of regular front ports. This means that the recirculation port limits the performance excessively.
@@
Our recirculation-based caching design avoids the potential bottleneck in the recirculation port as follows. First, the switch never recirculates requests. Second, only a small, constant number of cache packets are recirculated. The time to recirculate and process a cache packet is a few hundred nanoseconds like normal packets. Third, the switch can process multiple packets simultaneously through inter-stage parallelization, reducing queueing delay between cache packets.
@@
Trade-off. There is a trade-off of using recirculation to achieve variable-length caching. Although we are free from size limitations, we should limit the cache size. This is because, in OrbitCache, requests should wait until cache packets handle them. For a cache packet, the time to read a request is impacted by the other in-flight cache packets. This means that the number of in-flight cache packets in the recirculation port determines the latency for cache serving. Although the switch can process many cache packets simultaneously, the request may wait excessively until being served if there are too many cache packets. The sacrifice of cache size for variable-length caching is backed by the small cache effect. As described in Section 2.1, caching a small number of hot items is enough to balance skewed workloads [15, 27].
@@
  • The switch should adapt to key popularity changes.
Figure 2: OrbitCache architecture.
3 OrbitCache Design
3.1 OrbitCache Architecture
Figure 2 shows the overview of the OrbitCache architecture.
Switch data plane. The switch data plane in OrbitCache consists of several custom tables and modules as follows.
@@ 3.1 OrbitCache Architecture
  • Cloning module comprises a few match-action tables for cloning reply packets. Packet cloning is done by the PRE, a hardware module in the switch ASIC [5].
Meanwhile, the switch invokes the custom processing logic only for OrbitCache packets by referring to reserved L4 ports. We use UDP to handle messages for better latency like existing works [27, 29, 21] while using TCP for top-$k$ item reports in cache updates. For normal packets, the switch only applies the traditional packet forwarding logic.
Switch control plane. The controller in the switch control plane performs cache updates and switch configurations. It evicts the least popular keys and inserts new hot keys based on server-side periodic top-$k$ reports for uncached keys and the switch-side report for cached keys. Note that while the controller handles key insertion, value fetching is done via the switch data plane. It also handles switch configurations like the rule update for packet forwarding tables.
Clients and servers. Clients should specify the operation type, the item key to request, and the key hash. Storage servers run a server application that acts as a shim layer that translates OrbitCache messages to API calls for key-value stores and vice versa. The server returns replies for read and write requests like regular storage servers. However, in cases of write requests for cached items, the server returns a write reply with the item value as well to fetch the latest value.
Figure 3: OrbitCache packet format.